Operation Summary for April 9, 2010 - April 12, 2010 |
The Integrated Luminosity for the period from 4/5/10 to 4/12/10 was 69.51 pb-1. Friday April 9 The day shift began with Operations in shot setup, the Recycler stash (409.1E10), and the Antiproton Source stack (23.8E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI and MiniBooNE experiments taking beam. During the day shift, Operations established store 7733 at 9:15 AM with an average initial luminosity (AIL) of 359.9E30. During the evening shift, Operations disabled beam to MTest at 6:37 PM. Saturday April 10 The midnight shift began with Operations monitoring store 7733 (77.12E30) and the stash (364.4E10), with Pbar stacking (26.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7733 at 1:30 AM and put the TeV into shot setup at 1:49 AM. Operations established store 7734 at 2:55 AM with an AIL of 351.6E30. Operations established beam to MTest at 4:16 AM. During the day shift, Operations monitored store 7734. During the evening shift, Operations disabled the MTest beam at 6:24 PM. Operations terminated store 7734 at 7:20 PM and put the TeV into shot setup and established store 7735 at 9:00 PM with an AIL of 365E30. Sunday April 11 The midnight shift began with Operations monitoring store 7735 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, a Linac quadrupole power supply tripped off and caused a short downtime; it was reset. Operations established beam to MTest at 4:36 AM. During the day shift, Operations terminated store 7735 at 1:40 PM and put the TeV into shot setup and established store 7736 at 3:00 PM with an AIL of 373.8E30. LRF1, LRF2 and LRF5 tripped causing minimal downtime; all were reset. During the evening shift, Operations disabled beam to MTest at 6:01PM. Monday April 12 The midnight shift began with Operations monitoring store 7736 and the stash, with Pbar stacking, with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7736 at 7:45 AM and put the TeV into shot setup. A minor downtime occurred when the I- vacuum valve closed; the valve was re-opened.
The Plans for Monday The plan for today is to put in a Tev store, then continue to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 04/12/2010 email Fermilab |
| Security, Privacy, Legal | |